When 0.40 g of impure zinc reacted with excess hydrochloric acid, 127mL of hydrogen gas were collected over water at 10 degrees C at a total pressure of 737.77 Torr. The vapor pressure of water at 10 degrees C is 9.21 Torr.

The reaction in question is: Zn (s) + 2HCl (aq) --> ZnCl2 (aq) + H2 (aq)

A.) what amount (in grams) of H2 gas was collected?

B.) what is the percentage of purity of the zinc, assuming that all of the zinc present reacted completely with HCl and that the impurities did not react with HCl to produce hydrogen gas?
Answer #1

Pressure of hydrogen gas = pressure of collection - vapor pressure = 737.77 - 9.21 = 728.56 torr = 728.56 / 760 atm

Moles of hydrogen (n) = PV / RT, where p is pressure in atm, V is volume in liters, R is gas constant and T is temperature

n = (728.56 / 760) * 0.127 / (0.0821 * 283) = 5.24 * 10-3 moles

a) mass of H2 = moles * molar mass = 5.24 * 10-3 * 2 = 1.05 * 10-2 g

b) From the equation, moles of H2 = moles of Zn reacted = 5.24 * 10-3

Mass of Zn = moles * molar mass = 5.24 * 10-3 * 65.4 = 0.343 g

% purity = mass of pure Zn / mass of sample * 100 = 0.343 * 100 / 0.4 = 85.75%